# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v1

@@ 3.2 Task-Agnostic Metrics
    computation of various time series from the raw data, such as the angle between 2 consecutive points, the distance between two points, and multiple derivations (see Section 3.2.1).
@@
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. The segmentation is described in Section 3.2.2
@@
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent metrics, described in Section 3.2.3.
@@
3.2.1 Time series
For every point $i\in[[2,N]]$ of a given clean data file $\mathcal{F}$, we compute:
  • The angle between 2 consecutive points $\alpha_{i}=arctan(y_{i}-y_{i-1},x_{i}-x_{i-1})$.
  • The $L2$ norm between 2 consecutive points $n_{i}=\sqrt{(x_{i}-x_{i-1})^{2}+(y_{i}-y_{i-1})^{2}}$.
Subsequently, we calculate the first and second derivatives for the following:
@@ 3.2.2 Segmentation
3.2.3 Time-Independent Metrics
From the time series computed previously, we extract a range of statistics to characterize the dynamic properties of the signal. The list of statistics includes, for a given signal $S=(s_{i})_{1\leq i\leq N}$:
  • Mean ($\mu):\frac{1}{N}\sum_{i=1}^{N}s_{i}$
  • Standard deviation ($\sigma):\sqrt{\frac{1}{N}\sum_{i=1}^{N}(s_{i}-\mu)^{2}}$
  • Amplitude ($A):max(S)-min(S)$
@@ 3.2.3 Time-Independent Metrics
3.3 Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks’ categories proposed in Section 2.2, with the addition of a section specific to the Numbers task.
3.3.1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor skills of participants and are mostly used for PD assessment. Most dynamic metrics have already been explored as task-agnostic metrics, this is why we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen’s position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
@@ 3.3.1 Point Tasks Metrics
    Average Radius measures the dispersion of the pen positions related to the geometric center, a first measurement of stability, which measure the amplitude of the variations.
@@
    Standard deviation of the Radius measures the variations in the distance from the center, a second measure more focus on intermittent tremors, not on regular vibrations.
@@
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all muscles, if only certain muscles of the arm were affected, we
@@
3.3.2 Spiral Tasks Metrics
@@
3.3.3 Numbers Task Metrics
@@
The average width of each answer quantifies the size of the handwriting. It is computed by subtracting the minimum $X$ coordinate from the maximum $X$ coordinate for each answer and then averaging these values. People with PD have been found to have a lower control of their movement amplitudes, resulting in longer strokes and larger handwriting [32].
@@
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line.
@@
3.3.4 Writing Tasks Metrics
@@
The pressure standard deviation metric is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this metric.
@@ 3.4 Statistical Analysis
3.5 Additional comparisons
The tasks analyzed consist of a carefully selected set of closely related tasks to enable comparative analysis of a specific type of task. Two types of comparisons are proposed:
  • Grouped metrics: Instead of measuring the significance of the metrics task by task, we compare metrics for groups of tasks.
@@ 3.5 Additional comparisons
3.5.1 Grouped metrics
We supposed that certain of the metrics extracted for a given type of task should have a coherent behavior across all tasks. To verify this, we grouped them: for each group of tasks, the same statistical comparisons as the ones detailed in section 3.4 are performed, but considering all recordings of a group of tasks instead of the recordings of a specific task. We measured the significance of all metrics for each of those task groups as one task:
@@ 3.5.1 Grouped metrics
Table 2 summarizes the tasks at hand sorted per group and the specific metrics proposed per group of tasks.
Table 2: summary of the tasks at hand and their specific metrics.
Group of tasks Tasks Description Specific Metrics
Point Right Maintain the pen vertically above a point without the Radius from centroid
Points Point Left moving or touching the tablet, using their right hand, Angle from Centroid
Point Sustained left hand, or dominant hand while sustaining the vowel /a/. Std of the Radius
@@
Drawing a round spiral from the center Ellipticity
Spirals Spiral Left using their right hand, left hand, Number of loops
or dominant hand while repeating PaTaKa. Distance between loops
@@
Answer 8 consecutive arithmetic Avg width of answer
Numbers - questions and writing the Change in Avg pressure
answer while saying it. Time per answer
@@
Copy Text Copy a small paragraph being shown, Pressure std.
Writing Copy Text Memory then copy the same from memory, Avg time between words
Freewrite finally writing a freeform paragraph. Avg distance between words
@@
Copy Image Copy an image being shown,
Drawing Copy Image Memory then copy the same from memory, CLIP similarity image-image
Draw Clock drawing a clock CLIP similarity text-image
Copy Cube and drawing a cube.
3.5.2 Difference between metrics
To study the variations between tasks and how changes of behavior might be signs of NDs, we compared the metrics extracted between 2 tasks. For each session, if a given participant had successfully completed both tasks of a pair, we took the difference between 2 metrics, and perform the same statistical analysis as described in section 3.4 for this new ’coupled’ metric. The considered couples are the following:
@@ 3.5.2 Difference between metrics
3.6 Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Support Vector Machine (SVM), K-Nearest Neighbors (KNN), Random Forest (RF), Extreme Gradient Boosting (XGBoost), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^3] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm.
A Nested Cross-Validation (NCV) scheme is being applied in our experiments. NCV is a common approach that can be adopted for metric selection and parameter tuning to obtain reliable classification accuracy and prevent overfitting [34, 35]. We are splitting the data set into ten outer folds, and each fold is held out for testing while the remaining $k-1$ folds (9 folds) are merged to form the outer training set. Each outer training set is further split into stratified inner folds for inner training and validation. We identify the 30 best metrics (IFMs only) and hyperparameters in the inner folds via grid search. An Extremely Randomized Trees Classifier is used to perform metric selection. [^4] The documentation of the Python library used to implement the Extremely Randomized Trees Classifier can be found at https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.ExtraTreesClassifier.html. It is a type of ensemble learning technique which aggregates the results of multiple de-correlated decision trees to output its classification result. Ten sets of unique best configurations are obtained after getting the best hyperparameter configuration for each outer iteration using their corresponding inner folds.
To improve stability in metric and hyperparameter selection with small datasets, we adapt the NCV process. We determine the final hyperparameter configuration by averaging (or using the mode of) the best configurations from ten iterations. The model is then trained and tested from scratch using this configuration on distinct outer training/test folds, ensuring no overlap between training and testing participants. This approach includes 10-fold cross-validation for hyperparameter tuning with all training data and applies NCV separately for each task. Data is standardized within each outer iteration using the training set’s mean and standard deviation. Figure 9 outlines the classification pipeline and NCV scheme.
@@ 4 Results
4.1 Task-agnostic Metrics Results
@@
Out of 54 task-agnostic metrics computed, 53 are significant ($p\leq 0.05$) at least for one task and one comparison between groups. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant metrics. We present the metrics most effective in distinguishing AD and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight metrics that exhibited robust behavior across multiple comparisons.
4.1.1 AD vs CTL
@@
Over the 54 task-agnostic metrics, 48 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 17 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table 3 across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 10.
@@
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one’s handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
@@
When comparing the tasks Copy Text and Copy Read Text, the difference in speed and horizontal acceleration is significantly higher for AD participants than CTL participants ($p$-value$/leq$0.001 for both), which shows how the multi-tasking has a way higher effect on cognitively participants.
@@
Table 3: $p$-values and AUC for the best 10 metrics for AD vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD vs CTL Copy Image Copy Text Draw Clock Freewrite
p-val AUC p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.006 0.700 $<$0.001 0.791 $<$0.001 0.761 $<$0.001 0.760
in-air shannon entropy $<$0.001 0.873 $<$0.001 0.887 $<$0.001 0.866 $<$0.001 0.769
in-air std angular acc. $<$0.001 0.737 $<$0.001 0.818 $<$0.001 0.768 $<$0.001 0.821
in-air std angular speed $<$0.001 0.730 $<$0.001 0.826 $<$0.001 0.799 $<$0.001 0.774
on-tablet avg. hor. acc. $<$0.001 0.732 0.205 0.605 $<$0.001 0.788 0.003 0.698
AD vs CTL Numbers Point Dominant Point Non-Dominant Point Sustained
p-val AUC p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.025 0.691 $<$0.001 0.726 0.077 0.627 $<$0.001 0.740
in-air avg. acc. 0.685 0.536 $<$0.001 0.792 $<$0.001 0.850 $<$0.001 0.809
in-air avg. hor. acc. 0.357 0.593 0.021 0.663 0.006 0.688 0.007 0.698
in-air avg. speed 0.922 0.507 $<$0.001 0.797 $<$0.001 0.857 $<$0.001 0.827
in-air shannon entropy $<$0.001 0.794 $<$0.001 0.865 $<$0.001 0.810 $<$0.001 0.872
in-air std angular acc. 0.014 0.727 $<$0.001 0.804 0.003 0.708 $<$0.001 0.789
in-air std acc. 0.346 0.602 $<$0.001 0.716 $<$0.001 0.766 0.045 0.655
in-air std angular speed 0.025 0.694 0.021 0.665 0.131 0.608 0.112 0.615
in-air std speed 0.357 0.593 0.025 0.650 $<$0.001 0.737 0.052 0.647
@@
4.1.2 PD vs CTL
Over the 54 task-agnostic metrics, 19 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 4 have p-values under 0.001. Table 4 shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
Table 4: $p$-values and AUC for the best 10 metrics for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.

### v2

@@ III-B Task-Agnostic Metrics
    computation of time series from the raw data, such as the angle between 2 consecutive points, the distance between two points, and multiple derivations (see Section III-B1).
@@
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. The segmentation is described in Section III-B2
@@
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent metrics, described in Section III-B3.
@@
III-B1 Time series
For every point of a given clean data file, we compute the angle and the $L_{2}$ norm between 2 consecutive points. Subsequently, we calculate the first and second derivatives for the following:
@@ III-B2 Segmentation
III-B3 Time-Independent Metrics
From the time series computed previously, we extract a range of statistics to characterize the dynamic properties of the signal, namely: the mean, the standard deviation and the amplitude of the signal.
@@ III-B3 Time-Independent Metrics
III-C Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks’ categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
III-C1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic metrics have already been explored as task-agnostic metrics, so we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen’s position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
@@ III-C1 Point Tasks Metrics
    Average Radius measures the dispersion of pen positions relative to the geometric center, providing a first measurement of stability that assesses the amplitude of variations for essential tremor.
@@
    Standard deviation of the Radius measures the variations in the distance from the center, a second measure more focus on dystonic tremors, for irregular vibrations.
@@
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all limb regions equally. An average angle ¿ 0 might implicate a strong tremor axis.
@@
III-C2 Spiral Tasks Metrics
@@
III-C3 Numbers Task Metrics
@@
The average width of each answer quantifies the size of the handwriting. It is computed by subtracting the minimum $X$ coordinate from the maximum $X$ coordinate for each answer and then averaging these values. People with PD have been found to have less control of their movement amplitudes and variance, resulting in compensatory longer strokes and larger handwriting under medication [40], as they are during the studied tasks.
@@
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line. This metric might help pick up dystonia features, involuntary muscle contractions commonly seen in neurodegenerative diseases.
@@
III-C4 Writing Tasks Metrics
@@
The pressure standard deviation metric is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this metric, as no pressure can be computed if the pen is not touching the tablet.
@@ III-D Statistical Analysis
III-E Additional comparisons
The analyzed tasks consist of a carefully selected subset of related activities, allowing for a comparative analysis of a specific task group. Two types of comparisons are proposed:
  • Grouped metrics: Instead of measuring the significance of the metrics individually, we compare metrics for groups of tasks, such as the average speed, for all spirals.
@@ III-E Additional comparisons
III-E1 Grouped metrics
We hypothesized that certain metrics extracted for a specific type of task would exhibit consistent behavior across all tasks. To test this hypothesis, we grouped the tasks together. For each group, we conducted the same statistical comparisons outlined in section III-D. However, instead of analyzing the recordings from individual tasks, we considered all recordings within each task group. We then measured the significance of all metrics for each of these task groups as if they were a single task.
@@ III-E1 Grouped metrics
Table II summarizes the tasks at hand sorted per group and the specific metrics proposed per group of tasks.
TABLE II: summary of the tasks at hand and their specific metrics.
Group Tasks Specific Metrics
Point Right Radius from centroid
Points Point Left Angle from Centroid
Point Sustained Std of the Radius
@@
Ellipticity
Spirals Spiral Left Number of loops
Distance between loops
@@
Avg width of answer
Numbers Numbers Change in Avg pressure
Time per answer
@@
Copy Text Pressure std.
Writing Copy Text Memory Avg time between words
Freewrite Avg distance between words
@@
Copy Image
Drawing Copy Image Memory CLIP similarity image-image
Draw Clock CLIP similarity text-image
Copy Cube
III-E2 Difference between metrics
To study the variations between tasks and how changes of behavior might be signs of NDs, we compared the metrics extracted between 2 tasks. For each session, if a given participant had successfully completed both tasks of a pair, we took the difference between 2 metrics, and perform the same statistical analysis as described in section III-D for this new ’coupled’ metric. The considered couples are the following:
@@ III-E2 Difference between metrics
III-F Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. Namely, it involves a binary classification (e.g., AD or CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the metrics’ effectiveness in distinguishing between the categories under investigation.
These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Support Vector Machine (SVM), K-Nearest Neighbors (KNN), Random Forest (RF), Extreme Gradient Boosting (XGBoost), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm.
Nested cross-validation (NCV) is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [42, 43]. NCV is structured around two sequential processes:
  1. 1.
    Outer loop: The data is divided into folds, where one fold serves as a test set, and the remaining folds form the training set. This is repeated so that each fold acts as the test set once, yielding a balanced estimate of model accuracy.
  2. 2.
    Inner loop: Within each outer training set, the data undergoes another round of splitting to optimize hyperparameters and select features. This tuning is isolated from the test set, ensuring the performance evaluation remains unbiased.
This separation of tuning and testing phases helps NCV deliver a realistic and robust measure of the model’s performance, making it especially effective for tasks where data scarcity could lead to overfitting.
@@ IV Results
IV-A Task-agnostic Metrics Results
@@
Out of 54 task-agnostic metrics computed, we found 53 were significant ($p\leq 0.05$) for, at least one task and one groups comparison. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant metrics. We present the metrics most effectively distinguishing AD and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight metrics that exhibited robust behavior across multiple comparisons.
IV-A1 AD vs CTL
@@
Over the 54 task-agnostic metrics, 48 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 17 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table III across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 10.
@@
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one’s handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
@@
When comparing the tasks Copy Text and Copy Read Text, the difference in speed and horizontal acceleration is significantly higher for AD participants than CTL participants ($p$-value $\leq$ 0.001 for both), which shows how the multi-tasking has a way higher effect on cognitive participants.
A limitation of our approach here is the lower number of participant for the AD/MCI category, which did not allowed us a separated analysis of the cognitive impairments against the participants presenting only AD’s signs.
@@
TABLE III: $p$-values and AUC for the best 10 metrics for AD vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD vs CTL Copy Image Copy Text Copy Read Text
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.006 0.700 $<$0.001 0.791 0.010 0.672
in-air shannon entropy $<$0.001 0.873 $<$0.001 0.887 $<$0.001 0.827
in-air std angular acc. $<$0.001 0.737 $<$0.001 0.818 0.003 0.702
in-air std angular speed $<$0.001 0.730 $<$0.001 0.826 0.005 0.692
on-tablet duration $<$0.001 0.801 $<$0.001 0.738 $<$0.001 0.751
on-tablet avg. speed 0.007 0.693 0.185 0.614 $<$0.001 0.784
AD vs CTL Numbers Draw Clock Freewrite
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.025 0.691 $<$0.001 0.761 $<$0.001 0.760
in-air shannon entropy $<$0.001 0.794 $<$0.001 0.866 $<$0.001 0.769
in-air std angular acc. 0.014 0.727 $<$0.001 0.768 $<$0.001 0.821
in-air std angular speed 0.025 0.694 $<$0.001 0.799 $<$0.001 0.774
on-tablet duration 0.036 0.681 $<$0.001 0.740 0.148 0.599
on-tablet avg. speed 0.346 0.602 $<$0.001 0.787 0.003 0.698
AD vs CTL Point Dom. Point Non Dom. Point Sustained
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. $<$0.001 0.726 0.077 0.627 $<$0.001 0.740
in-air avg. acc. $<$0.001 0.792 $<$0.001 0.850 $<$0.001 0.809
in-air avg. speed $<$0.001 0.797 $<$0.001 0.857 $<$0.001 0.827
in-air shannon entropy $<$0.001 0.865 $<$0.001 0.810 $<$0.001 0.872
in-air std angular acc. $<$0.001 0.804 0.003 0.708 $<$0.001 0.789
in-air std angular speed 0.021 0.665 0.131 0.608 0.112 0.615
@@
IV-A2 PD vs CTL
Over the 54 task-agnostic metrics, 19 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 4 have p-values under 0.001. Table IV shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
TABLE IV: $p$-values and AUC for the best 10 metrics for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.